OpenStudy (anonymous):

0 = x^+8x-65 ?? how to solve using factoring and quadratic formula ???

OpenStudy (anonymous):

Quadratic formula is b+/- \[\sqrt{-b-4ac}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[x^2+8x-65=0\] \[(x-5)(x+13)=0\] \[x=5\] or \[x=-13\]

OpenStudy (anonymous):

easier to factor than using th formula if you can

OpenStudy (anonymous):

did she post the ^2 ?

OpenStudy (anonymous):

okay thank you